A silver wire has a cross sectional area A = 2.0 mm2. A total of 9.4 × 1018 electrons pass through the wire in 3.0 s. The conduction electron density in silver is 5.8 × 1028 electrons/m3 and e = 1.60 × 10-19 C. What is the drift velocity of these electrons?
Answer #1

here,

area , A = 2 mm^2 = 2 * 10^-6 m^2

number of electrons , n = 9.4 * 10^18 electrons

time , t = 3 s

current in the wire , I = n * e /t

I = 9.4 * 10^18 * 1.6 * 10^-19 /( 3)

I = 0.501 A

The conduction electron density in silver, N = 5.8 × 10^28 electrons/m3

the drift velocity of these electrons , v = I/( N * a * e)

v = 0.501 /( 5.8 * 10^28 * 2 * 10^-6 * 1.6 * 10^-19)

v = 2.7 * 10^-5 m/s

the drift speed of electrons is 2.7 * 10^-5 m/s